OpenStudy (anonymous):

which equation would provide the best estimate of 32% of 44.7? i will meadel and fan 1/3 x 4=n 2/3x45=n 1/3x 45=n 2/3x 44=n

OpenStudy (lieutenantgeneral):

which equation would provide the best estimate of 32% of 44.7? = 44.7 * .32 14.304 32% = 8/25 = 1/3 * 45 or 2/3 * 45